1) Look Manufacturing Corporation has a traditional costing system in which it applies manufacturing overhead to its products using a predetermined overhead rate based on direct labor-hours (DLHs). The company has two products, N06D and M09K, about which it has provided the following data:

N06D M09K
Direct materials per unit $ 30.80 $ 64.10
Direct labor per unit $ 6.00 $ 28.00
Direct labor-hours per unit 0.20 1.00
Annual production (units) 48,800 21,300

The company's estimated total manufacturing overhead for the year is $1,866,884 and the company's estimated total direct labor-hours for the year is 31,060.

The company is considering using a variation of activity-based costing to determine its unit product costs for external reports. Data for this proposed activity-based costing system appear below:

Activities and Activity Measures Estimated
Overhead Cost
Supporting direct labor (DLHs) $ 1,024,980
Setting up machines (setups) 419,648
Parts administration (part types) 422,256
Total $ 1,866,884
Expected Activity
N06D M09K Total
DLHs 9,760 21,300 31,060
Setups 1,580 948 2,528
Part types 651 261 912

The manufacturing overhead that would be applied to a unit of product M09K under the activity-based costing system is closest to:

2) Adelberg Corporation makes two products: Product A and Product B. Annual production and sales are 3,000 units of Product A and 375 units of Product B. The company has traditionally used direct labor-hours as the basis for applying all manufacturing overhead to products. Product A requires 0.2 direct labor-hours per unit and Product B requires 0.8 direct labor-hours per unit. The total estimated overhead for next period is $100,720.

The company is considering switching to an activity-based costing system for the purpose of computing unit product costs for external reports. The new activity-based costing system would have three overhead activity cost pools--Activity 1, Activity 2, and General Factory--with estimated overhead costs and expected activity as follows:

Expected Activity
Activity Cost Pool Estimated Overhead Costs Product A Product B Total
Activity 1 $ 40,600 900 500 1,400
Activity 2 28,350 1,000 500 1,500
General Factory 31,770 600 300 900
Total $ 100,720

(Note: The General Factory activity cost pool's costs are allocated on the basis of direct labor-hours.)

The overhead cost per unit of Product B under the activity-based costing system is closest to:
Answer #1

Total manufacturing overhead applied to M09K = (Estimated overhead cost/Total Activity) *Activity Consumed

= (1024980*21300/31060 + 419648*948/2528 + 422256*261/912)

= $981,111

Cost per unit = 981,111/21300

= $46.0615 per unit

2)Total overhead assigned = (40600*500/1400 + 28350*500/1500 + 31770*300/900)

= $34,540

Per unit = 34,540/375

= $92.1067